Assume that a coin is flipped where the probability of coin lands Heads is 0.49. The coin is flipped once more. This time,
Answer #1

Solution Given P(H) = 0.49 P/H1H) = 0.38 P(TIH) = (-0.38 P(TIH) = 0.62 Now the values of x,y along with their probabilities f= Now to find E(y) E(Y) = 2 x (0.1862 + 0.1938) 2 x 0.38 E(4) 0.76 Now to find E(x²) E(X) 2 x 0.186 2 + 1 2 (0.3 0. 3038 + 0.Now to find Varly) Var(y) = E(+2) - [[(y)] ² Varly) = 1.52-(0.76² = 1.52 -0.5776 Var (4) 0.9424 Now to find Elxy) E(XY) = 2x2Now 4 to find the correlation between X andy Cor (x, y) pa var(x) Var(y) -0.0094 6.3799 (0.9924) -0.0094 = P 0.3770 -0.0094 으